# CCFExp: Facial Image Synthesis with Cycle Cross-Fusion Diffusion Model for Facial Paralysis Individuals

### v1

@@ 4.1 Datasets
The Annotated Facial Landmarks for Facial Palsy (AFLFP) [64] dataset is a comprehensive and laboratory-controlled collection designed to support the analysis of facial paralysis. It consists of facial images from 88 subjects, each participating in 16 different expression videos, such as brow raise, closed smile, gentle eye closure, and open smile. For each expression, the dataset captures keyframes representing four critical states: neutral, onset, a mid-state between onset and peak, and peak. In total, the AFLFP dataset contains 5,632 facial images, with 1,408 images corresponding to each key state and 64 images per subject. Each image is meticulously annotated with 68 facial landmarks, providing detailed geometric information crucial for modeling facial movements and asymmetries associated with facial paralysis.
@@
The Massachusetts Eye and Ear Infirmary (MEEI) [17] dataset is a carefully curated collection focused on facial palsy, comprising both videos and high-resolution images. The dataset includes 60 videos from a diverse group of participants, consisting of 9 healthy subjects and 51 patients with varying degrees of facial paralysis. Each video captures 8 distinct facial movements, which are also provided as individual images, resulting in a total of 480 high-resolution images. The subjects in the MEEI dataset are categorized based on their eFACE scores, which assess the severity of facial palsy. The categories include normal (96–100), near-normal (91–95), mild (80–90), moderate (70–79), severe (60–69), and complete flaccid or nonflaccid facial palsy (below 60). This classification enables detailed analysis of facial movement across a broad spectrum of facial palsy severity.
@@
4.2 Evaluation Metrics
Fréchet Inception Distance (FID) [18] is widely used to evaluate the quality of image generation. It measures the distance between the distribution of generated images and real images by comparing the statistics of feature representations extracted from a pre-trained Inception network. A lower FID score indicates that the generated images are more similar to the real images, both in terms of visual quality and diversity. However, some studies [7, 45] have shown that inconsistencies in the class distribution between the generated dataset and the real dataset can affect the calculation of the FID score. To address this issue, we propose the aligned FID (aFID) to evaluate the quality of the multi-condition diffusion model. Specifically, for each facial palsy patient in the test set, we first use images of other individuals as identity images and the patient’s condition image as the facial palsy expression image. This process synthesizes an image where the patient’s facial palsy expression is transferred onto another individual. We then use this generated dataset as the facial palsy expression images and the original facial palsy patient’s images as identity images to generate a second set of synthetic images. Finally, we calculate the FID score between this generated dataset and the real dataset. Ideally, the images in these two datasets should be identical.
Figure 5: Visualization of ablation results. CCFExp is the standard version, CCFExp1 is the CCFExp without facial landmarks, CCFExp2 excludes the cycle training strategy and CCFExp3 operates in the absence of the cross-fusion module.
@@ 4.3 Implementation details
CCFExp - $\times$ $\checkmark$ $\checkmark$ 50.47 17.31 16.05 0.45 0.40 0.31 0.33 0.25 0.28
@@
For the identity feature extractor $E_{ID}$, we utilized the same IR-50 [9] as DCFace [30]. Intermediate features from this network were injected into the cross-fusion module, and the final features were incorporated into the model prediction network with timesteps, and after initialization from DCFace, the model is trained end-to-end. For both the identity model and the face model within the expression extracting module, we also used the IR-50 architecture. After the same initialization, the ID model’s weights were fixed, and the module was trained on the AffectNet [41] dataset, after which the module’s weights were all fixed. For the landmark extractor $E_{LM}$ and detector $D_{LM}$, we used the MobileFaceNet [6] and fixed its weights to obtain the landmark information. For each patient’s video in the MEEI dataset, we performed frame extraction. Keyframes representing eight different expressions for each facial palsy patient were manually selected. Each expression consists of 32 frames, capturing the full expression process from onset, development, peak, dissipation, to a neutral state. For the images in AFLFP and MEEI, we first performed face extraction and segmented them into 112x112 facial images. After removing low-quality images and those with poor segmentation quality, we ultimately obtained 91,168 facial images from 313 individuals. We used 88% of the data (80,288 images from 276 individuals) for the model’s training set, while 12% of the data (10,880 images from 37 individuals) was set aside as the test set for subsequent evaluation metric calculations. CCFExp was trained on the training set using the AdamW Optimizer [33, 37] for 10 epochs with a learning rate of 0.0001, utilizing two NVIDIA RTX 4090 GPUs.
@@ 4.4 Ablation Study
To evaluate the contribution of each component in our model, we conducted a series of ablation studies. These experiments assess the impact of removing or altering key modules on the model’s performance. The visualization results of the ablation study are shown in Fig 5. To evaluate the effectiveness of image generation from different perspectives, for the synthesized images, in addition to aFID, we also compute Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM) [63], Learned Perceptual Image Patch Similarity (LPIPS) [69], Perceptual Loss (PL) [25], on unseen data, by comparing them separately with the identity image XID and the facial palsy expression image XStyle. The quantitative results are shown in Tab. 1.
@@
To investigate the role of facial landmark features in image generation, we remove the facial landmark condition from CCFExp. The model no longer learns landmark features, and in the fusion module, only identity features and expression features are used for information exchange through query key interactions and concatenation. During training, the model also stops using the landmark loss for weight updates. The results in Tab. 1 show a significant increase in aFID, and the pixel-level comparison metrics for facial palsy expression also exhibit a clear decline. As seen in Fig. 5, the model is largely unable to correctly learn facial palsy expressions. This highlights the importance of facial landmark features in synthesizing detailed facial palsy expressions.
@@ 4.5 Comparison with Previous Methods
We compared CCFExp with a similar facial synthesis task, DCFace [30], using 3x3 and 5x5 grid sizes on unseen images. The comparison results are shown in the upper part of Tab. 1. Although DCFace 5x5 achieves better scores in several generated data and facial palsy expression data metrics, this does not necessarily indicate a higher degree of facial palsy feature transfer. Pixel-level and perception-based metrics not only focus on facial expressions but also consider factors like lighting and background, which are not primary considerations in the task of facial palsy synthesis. In terms of aFID and identity image comparison metrics, CCFExp significantly outperforms DCFace, reaching the state-of-the-art performance.
@@ 5 Conclusion
Our research proposes Cycle Cross-Fusion Expression Generative Model (CCFExp) , based on the conditional diffusion model, utilizing multiple feature extractors to accurately capture various facial features, employing a cross-fusion module to integrate these features and applying a cycle training strategy to improve training efficiency and effectiveness. As a result, CCFExp generates realistic facial palsy expression images, providing valuable training data for machine learning systems used in facial palsy diagnosis and treatment assistance. Results in Tab. 1 indicates that CCFExp surpasses existing methods in several metrics, including aFID, PSNR, and SSIM, achieving excellent performance.
Limitations
Although CCFExp has shown promising results in the field of facial palsy feature transfer, a large amount of facial palsy expression data, especially manually annotated datasets, is still needed to further enhance the model’s performance and generalization capability. We encourage more researchers to participate in the collection of facial palsy data, which will help accelerate the early application of large-scale deep learning models in this field.

### v3

@@ 4.1 Datasets
The Annotated Facial Landmarks for Facial Palsy (AFLFP) [65] dataset is a comprehensive and laboratory-controlled collection designed to support the analysis of facial paralysis. It consists of facial images from 88 subjects, each participating in 16 different expression videos, such as brow raise, closed smile, gentle eye closure, and open smile. For each expression, the dataset captures keyframes representing four critical states: neutral, onset, a mid-state between onset and peak, and peak. In total, the AFLFP dataset contains 5,632 facial images, with 1,408 images corresponding to each key state and 64 images per subject.
@@
The Massachusetts Eye and Ear Infirmary (MEEI) [18] dataset is a carefully curated collection focused on facial palsy, comprising both videos and high-resolution images. The dataset includes 60 videos from a diverse group of participants, consisting of 9 healthy subjects and 51 patients with varying degrees of facial paralysis. Each video captures 8 distinct facial movements, which are also provided as individual images, resulting in a total of 480 high-resolution images. The subjects in the MEEI dataset are categorized based on their eFACE [3] scores, which assess the severity of facial palsy.
4.2 Evaluation Metrics
Figure 4: A schematic diagram of the aFID calculation method. Calculate the FID value between the twice-synthesized data and the original data.
To evaluate the effectiveness of image generation from different perspectives, for the synthesized images, we compute Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM) [64], Learned Perceptual Image Patch Similarity (LPIPS) [70], Perceptual Loss (PL) [26], on unseen data. For this multi-condition synthesis task, we compute these metrics separately for identity images and facial paralysis expression images to simultaneously evaluate identity preservation and expression transfer.
Fréchet Inception Distance (FID) [19] is widely used to evaluate the quality of image generation. It measures the distance between the distribution of generated images and real images by comparing the statistics of feature representations extracted from a pre-trained Inception network. A lower FID score indicates that the generated images are more similar to the real images, both in terms of visual quality and diversity. However, some studies [8, 46] have shown that inconsistencies in the class distribution between the generated dataset and the real dataset can affect the calculation of the FID score.
To address this issue, we propose the aligned FID (aFID) (See Fig. 4) to evaluate the quality of the multi-condition diffusion model. Specifically, for each facial palsy patient in the test set, we first use images of other individuals as identity images and the patient’s condition image as the facial palsy expression image. This process synthesizes an image where the patient’s facial palsy expression is transferred onto another individual. We then use this generated dataset as the facial palsy expression images and the original facial palsy patient’s images as identity images to generate a second set of synthetic images. Finally, we calculate the FID score between this generated dataset and the real dataset. Ideally, the images in these two datasets should be identical.
Figure 5: Visualized experimental results. CFCPalsy is the standard version, CFCPalsy1 is the CFCPalsy without facial landmarks, CFCPalsy2 excludes the cycle training strategy and CFCPalsy3 operates in the absence of the cross-fusion module.
@@ 4.3 Implementation details
CFCPalsy - $\times$ $\checkmark$ $\checkmark$ 50.47 17.31 16.05 0.45 0.40 0.31 0.33 0.25 0.28
@@
For the identity feature extractor $E_{ID}$, we utilized the same IR-50 [10] as DCFace [31]. Intermediate features from this network were injected into the cross-fusion module, and the final features were incorporated into the model prediction network with timesteps, and after initialization from DCFace, the model is trained end-to-end. For both the identity model and the face model within the expression extracting module, we also used the IR-50 architecture. After the same initialization, the ID model’s weights were fixed, and the module was trained on the AffectNet [42] dataset, after which the module’s weights were all fixed. For the landmark extractor $E_{LM}$ and detector $D_{LM}$, we used the MobileFaceNet [7] and fixed its weights to obtain the landmark information. For each patient’s video in the MEEI dataset, we performed frame extraction. Keyframes representing eight different expressions for each facial palsy patient were manually selected. Each expression consists of 32 frames, capturing the full expression process from onset, development, peak, dissipation, to a neutral state. For the images in AFLFP and MEEI, we first performed face extraction and segmented them into 112x112 facial images. After removing low-quality images and those with poor segmentation quality, we ultimately obtained 91,168 facial images from 313 individuals. We used 88% of the data (80,288 images from 276 individuals) for the model’s training set, while 12% of the data (10,880 images from 37 individuals) was set aside as the test set for subsequent evaluation metric calculations. CFCPalsy was trained on the training set using the AdamW Optimizer [34, 38] for 10 epochs with a learning rate of 0.0001, utilizing two NVIDIA RTX 4090 GPUs.
@@ 4.4 Ablation Study
To evaluate the contribution of each component in our model, we conducted a series of ablation studies. These experiments assess the impact of removing or altering key modules on the model’s performance. The visualization results of the ablation study are shown in Fig 5. The quantitative results are shown in Tab. 1.
@@
To investigate the role of facial landmark features in image generation, we remove the facial landmark condition from CFCPalsy. The model no longer learns landmark features, and in the fusion module, only identity features and expression features are used for information exchange through query key interactions and concatenation. During training, the model also stops using the landmark loss for weight updates. The results in Tab. 1 show a significant increase in aFID, and the pixel-level comparison metrics for facial palsy expression also exhibit a clear decline. As seen in Fig. 5, the model is largely unable to correctly learn facial palsy expressions. This highlights the importance of facial landmark features in synthesizing detailed facial palsy expressions.
@@ 4.5 Comparison with Previous Methods
We compared CFCPalsy with a similar facial synthesis task, DCFace [31], using 3x3 and 5x5 grid sizes on unseen images. The visual comparison results and numerical comparison results are shown in Fig. 5 and Tab. 1, respectively. Although DCFace 5x5 achieves better scores in several generated data and facial palsy expression data metrics, this does not necessarily indicate a higher degree of facial palsy feature transfer. Pixel-level and perception-based metrics not only focus on facial expressions but also consider factors like lighting and background, which are not primary considerations in the task of facial palsy synthesis. Compared to DCFace 5x5, aFID of images generated by CFCPalsy decreases by approximately 22%, and CFCPalsy also demonstrates a clear advantage in terms of identity feature preservation.
@@ 5 Conclusion
Our research proposes Cross-Fusion Cycle Palsy Expression Generative Model (CFCPalsy) , based on the conditional diffusion model, utilizing multiple feature extractors to accurately capture various facial features, employing a cross-fusion module to integrate these features and applying a cycle training strategy to improve training efficiency and effectiveness. As a result, CFCPalsy generates realistic facial palsy expression images, providing valuable training data for machine learning systems used in facial palsy diagnosis and treatment assistance. Results indicates that CFCPalsy surpasses existing methods in several metrics, including aFID, PSNR, and SSIM, achieving excellent performance.